How many orders will be visible if the wavelength of the incident radiation is 4800 Å and the number of lines on the grating is 2500 lines per inch.


1
Expert's answer
2020-12-15T07:31:47-0500


dsinθ = mλ . when λ = 90o m = -mmax and m = mmax

where d is the diffraction grating line spacing, λ is the wavelength of the light source, θ is the

angle form the grating to the screen, and m = -mmax .... 0 ... mmax (integers)


We know that b + d  the maximum value of θ is 900, that is sinθ = 1. Therefore, the number of orders visible with the granting is given as b + d = nλ

the grating is 2500 lines per inch.

m = "\frac{b + d}{λ}"

here b + d = 2500 lines per inch to 2.54 / 2500 to per cm

and λ = 5000"\dot{A}" = 5000 * 10-8cm

m = "\frac{2.54}{2500 * 4800 * 10^{-8}}" = 21.6

m > 21

21 orders will be visible